In a random sample of 80 people, 44 consider themselves as basketball fans. Compute a 99% confidence interval for the true proportion of people consider themselves as basketball fans and fill in the blanks appropriately.

We are______ % confidence that the true proportion of people consider themselves as basketball fans is between ____and_____ . (Keep 3 decimal places)
Answer #1

Solution - Given that x=44 n=80 so, = x = 44 -0.55 1- ² = 1-0.55 = 0.45 is, At 99% confidence level the z 2= -99% 2 = -0.99=